# Robust Spatial Confounding Adjustment via Basis Voting

### v1

@@ 8 Acknowledgement
$\sigma^{2}_{j,d}=\nabla g^{\top}\bm{\Sigma}_{AB}\nabla g=$ $\dfrac{\sigma^{2}_{U,j}+\sigma^{2}_{\epsilon}}{{\alpha_{x}^{(j)}}^{2}}+\left(\dfrac{\alpha_{u}^{(j)}}{{\alpha_{x}^{(j)}}^{2}}\right)^{2}\sigma^{2}_{X,j}-2\dfrac{\alpha_{u}^{(j)}}{{\alpha_{x}^{(j)}}^{3}}\sigma_{U,X,j}$
@@
Thus, if $j\in O_{X}$, then $\alpha_{u}^{(j)}=0$ and thus $r_{j,d}=0$, which means $\sigma_{j,d}^{2}=\frac{\sigma^{2}_{U,j}+\sigma^{2}_{\epsilon}}{\alpha_{x}^{(j)^{2}}}$. Thus for such a $j\in O_{X}$, we have:
@@
This completes the proof of the first part lemma. Next, we move to proving the uniform tightness of a matrix deviation. Before that, we prove something very general:
Proposition 1.
[…]
where $||\cdot||_{2}$ is the spectral norm.
Proof.
$\bm{A}_{n}^{-1}=$ $(\bm{A}_{n}+\bm{A}-\bm{A})^{-1}$
$=$ $(\bm{A}+\bm{\Delta}_{n})^{-1}$
$=$ $\bm{A}^{-1}\left(\bm{I}+\bm{\Delta}_{n}\bm{A}^{-1}\right)^{-1}$
$=$ $\bm{A}^{-1}\sum_{k\geq 0}\left(-\bm{\Delta}_{n}\bm{A}^{-1}\right)^{k}$
[…]
This completes the proof of the proposition. ∎
We will be using this result in proving the second part of our lemma. Recall how We defined the matrix $\bm{A}_{j,d}$ as:
@@
Thus, we have:
$\begin{split}\hat{\beta}_{-j,n}-\beta=&\dfrac{\frac{1}{n}{\bm{X}}_{n}^{\top}\bm{P}_{d}(\bm{I}_{n}-\bm{P}_{-j,d})\bm{P}_{d}(\bm{U}_{n}+\bm{\epsilon}_{n})}{\frac{1}{n}{\bm{X}}_{n}^{\top}\bm{P}_{d}(\bm{I}_{n}-\bm{P}_{-j,d})\bm{P}_{d}\bm{X}_{n}}\\ =&\dfrac{\frac{1}{n}{\bm{X}}_{n}^{\top}\bm{H}_{d}\bm{A}_{j,d}\frac{1}{n}{\bm{H}}_{d}^{\top}(\bm{U}_{n}+\bm{\epsilon}_{n})}{\frac{1}{n}{\bm{X}}_{n}^{\top}\bm{H}_{d}\bm{A}_{j,d}\frac{1}{n}{\bm{H}}_{d}^{\top}\bm{X}_{n}}\\ \xrightarrow[n\to\infty]{P}&\frac{\bm{\alpha}_{x,d}^{\top}A^{lim}_{j,d}(\bm{\alpha}_{u,d}+\bm{0}_{d})}{{\bm{\alpha}}_{x,d}^{\top}A^{lim}_{j,d}\bm{\alpha}_{x,d}}=\frac{\alpha_{x}^{(j)}\alpha_{u}^{(j)}}{\alpha_{x}^{(j)^{2}}}=\frac{\alpha_{u}^{(j)}}{\alpha_{x}^{(j)}}=r_{j,d}\end{split}$ (22)
This completes the proof for the consistency of the bias of our drop-one estimator. Next, we will show the asymptotic distribution of the drop one estimator when the index $j$ of the dropped basis function lies in $O_{X}$, i.e. $\alpha_{u}^{(j)}=0$. Notice that $\bm{U}_{n}$ can be decomposed as $\bm{U}_{n}=\bm{U}_{n,d}+(1-\lambda_{d})\bm{U}_{n,>d}$, where the above condition also means that $\bm{U}_{n,d}\in\mathscr{C}(\bm{H}_{-j,d})$. Thus, we have:
@@
Using this, for $j\in O_{X}$, we can simplify the bias of the estimator as:
@@
Thus, we can say that $\sqrt{n}\left(\hat{\beta}_{-j,n}-\beta\right)$ has the same asymptotic distribution as that of the quantity $\sqrt{n}L_{n}=R_{n}\left(\bm{\alpha}_{x,d}^{\top}A^{lim}_{j,d}\bm{\alpha}_{x,d}\right)^{-1}$. So let us start deriving the asymptotic distribution of $R_{n}$. For that, first let us denote the vector $\bm{h}_{d,i}=(h_{1}(S_{i}),\cdots,h_{d}(S_{i}))^{\top}$. Hence we can write down $R_{n}$ as:
@@
Thus, using Multivariate CLT, we can say that:
@@
Finally, observe that when $U$ has a finite expansion with $d$ basis functions, i.e. when $\lambda_{d}=1$, then for $j\in O_{X}$ recalling that error vector $\bm{\epsilon}_{n}$ is independent of the sampled location vector $\bm{S}_{n}$, from Equation 24, we have:
@@
To prove this Lemma, we will first state and prove another one, which will be used in the proof.
Lemma S1.4.
Let $(\mathcal{S},\mathcal{F},f_{S})$ be a probability space and $\{h_{k}\}_{k\geq 1}\subset\mathcal{L}^{4}(\mathcal{D},f_{S})$ an orthonormal system. Fix $d\in\mathbb{N}$. Assume that for each $j\leq d$, the zero set $\{s:h_{j}(s)=0\}$ has $f_{S}$-measure 0 (equivalently, $h_{j}^{2}>0$ $f_{S}$-a.e.). For $j\leq d$, set $V_{j}\coloneqq\operatorname{span}\{h_{k}:\,1\leq k\leq d,\ k\neq j\}$ and define the set $\mathbb{S}_{j}\coloneqq\{v\in V_{j}:\ \|v\|_{f_{S}}=1\}$ and $Q_{j}(v)\coloneqq\int h_{j}^{2}\,v^{2}\,f_{S}$. Then
$c_{\ast}\coloneqq\inf_{1\leq j\leq d}\ \inf_{v\in\mathbb{S}_{j}}Q_{j}(v)>0.$
Proof.
$Q_{j}(v)=\int h_{j}^{2}v^{2}\,f_{S}>0.$
Thus $Q_{j}$ is strictly positive on $\mathbb{S}_{j}$.
Now we can prove Lemma 2, fix $1\leq j\leq d$ and assume $\alpha_{u}^{(j)}=0$ i.e. $j\in O_{X}$. Write $R_{j}\coloneqq\sum_{k\leq d,\,k\neq j}\alpha_{u}^{(j)}h_{k}$ and $T\coloneqq U_{>d}$. Set
$A_{j}\coloneqq\int h_{j}^{2}R_{j}^{2}\,f_{S}\,ds,\qquad B_{j}\coloneqq\int h_{j}^{2}T^{2}\,f_{S}\,ds,\qquad C_{j}\coloneqq\int h_{j}^{2}R_{j}T\,f_{S}\,ds.$
Then
By Cauchy-Schwarz, $|C_{j}|\leq\sqrt{A_{j}B_{j}}$, hence
Second, by Assumptions (i) and (iii),
Thus,

### v2

@@ 8 Acknowledgement
$\sigma^{2}_{j}=\nabla g^{\top}\bm{\Sigma}_{AB}\nabla g=$ $\dfrac{\sigma^{2}_{U,j}+\sigma^{2}_{\epsilon}}{{\alpha_{x}^{(j)}}^{2}}+\left(\dfrac{\alpha_{u}^{(j)}}{{\alpha_{x}^{(j)}}^{2}}\right)^{2}\sigma^{2}_{X,j}-2\dfrac{\alpha_{u}^{(j)}}{{\alpha_{x}^{(j)}}^{3}}\sigma_{U,X,j}$
@@
Thus, if $j\in O_{X}$, then $\alpha_{u}^{(j)}=0$ and thus $r_{j}=0$, which means $\sigma_{j}^{2}=\frac{\sigma^{2}_{U,j}+\sigma^{2}_{\epsilon}}{\alpha_{x}^{(j)^{2}}}$. Thus for such a $j\in O_{X}$, we have:
@@
This completes the proof of the first part lemma. Next, we move to proving the uniform tightness of a matrix deviation.
We will be using Lemma S3.3 in proving the second part of our lemma. Recall how We defined the matrix $\bm{A}_{j,d}$ as:
@@
Thus using the above statements and Lemma S4.7, we have:
$\begin{split}\hat{\beta}_{-j,n}-\beta=&\dfrac{\frac{1}{n}{\bm{X}}_{n}^{\top}\bm{P}_{d}(\bm{I}_{n}-\bm{P}_{-j,d})\bm{P}_{d}(\bm{U}_{n}+\bm{\epsilon}_{n})}{\frac{1}{n}{\bm{X}}_{n}^{\top}\bm{P}_{d}(\bm{I}_{n}-\bm{P}_{-j,d})\bm{P}_{d}\bm{X}_{n}}\\ =&\dfrac{\frac{1}{n}{\bm{X}}_{n}^{\top}\bm{H}_{d}\bm{A}_{j,d}\frac{1}{n}{\bm{H}}_{d}^{\top}(\bm{U}_{n}+\bm{\epsilon}_{n})}{\frac{1}{n}{\bm{X}}_{n}^{\top}\bm{H}_{d}\bm{A}_{j,d}\frac{1}{n}{\bm{H}}_{d}^{\top}\bm{X}_{n}}\\ \xrightarrow[n\to\infty]{P}&\frac{\bm{\alpha}_{x,d}^{\top}A^{lim}_{j,d}(\bm{\alpha}_{u,d}+\bm{0}_{d})}{{\bm{\alpha}}_{x,d}^{\top}A^{lim}_{j,d}\bm{\alpha}_{x,d}}=\frac{\alpha_{x}^{(j)}\alpha_{u}^{(j)}}{\alpha_{x}^{(j)^{2}}}=\frac{\alpha_{u}^{(j)}}{\alpha_{x}^{(j)}}=r_{j}\end{split}$ (22)
This completes the proof for the consistency of the bias of our drop-one estimator.
Next, we will show the asymptotic distribution of the drop one estimator when the index $j$ of the dropped basis function lies in $O_{X,d}$, i.e. $\alpha_{x}^{(j)}\neq 0$ and $\alpha_{u}^{(j)}=0$. Notice that $\bm{U}_{n}$ can be decomposed as $\bm{U}_{n}=\bm{U}_{n,d}+(1-\lambda_{d})\bm{U}_{n,>d}$, where the above condition also means that $\bm{U}_{n,d}\in\mathscr{C}(\bm{H}_{-j,d})$. Thus, we have:
@@
Using this, for $j\in O_{X,d}$, we can simplify the bias of the estimator as:
@@
Thus, we can say that $\sqrt{n}\left(\hat{\beta}_{-j,n}-\beta\right)$ has the same asymptotic distribution as that of the quantity $\sqrt{n}L_{n}=R_{n}\left(\bm{\alpha}_{x,d}^{\top}A^{lim}_{j,d}\bm{\alpha}_{x,d}\right)^{-1}$. So let us start deriving the asymptotic distribution of $R_{n}$:
@@
[…]
$\mathbb{E}h_{j}(S)U_{>d,k}(S)=\int h_{j}(s)X_{>d,k}(s)\,f_{S}ds=$ $\sum_{l=d+1}^{k}\alpha_{u}^{(l)}\int h_{j}(s)h_{l}(s)\,f_{S}\,ds$
$=$ $\sum_{l=d+1}^{k}\alpha_{u}^{(l)}\mathbb{I}(l=j)=0$
The reason for the last equality is that $l\geq d+1$ and $l\leq d$. Thus Dominated Convergence Theorem (DCT) implies that:
$\mathbb{E}h_{j}(S)U_{>d,k}(S)\xrightarrow[]{k\to\infty}\mathbb{E}h_{j}(S)U_{>d}(S)=\sum_{l=d+1}^{\infty}\alpha_{u}^{(l)}\mathbb{I}(l=j)=0$
This shows why the expectation of $h_{j}(S)U_{>d}(S)$ is 0. Thus, using Multivariate CLT, we can say that:
@@
Finally, observe that when $U$ has a finite expansion with $d$ basis functions, i.e. when $\lambda_{d}=1$, then for $j\in O_{X,d}$ recalling that error vector $\bm{\epsilon}_{n}$ is independent of the sampled location vector $\bm{S}_{n}$, from Equation 24, we have:
@@
Recall from Theorems 2 and 3, the asymptotic variance of the drop-one estimator for $j\in O_{X,d}$ is $avar\left(\hat{\beta}_{-j,n}\right)=\dfrac{(1-\lambda_{d})\sigma^{2}_{U,j,>d}+\sigma^{2}_{\epsilon}}{{\alpha_{x}^{(j)}}^{2}}$ and that for the projection estimator is $avar\left(\hat{\beta}^{PROJ}_{j,n}\right)=\frac{\sigma^{2}_{U,j}+\sigma^{2}_{\epsilon}}{{\alpha_{x}^{(j)}}^{2}}$. Let us compare the asymptotic variances under two scenarios:
Case 1: $(\lambda_{d}=1)$ This scenario implies that with respect to the basis function used for constructing the estimator, the unmeasured confounder assumes a finite expansion i.e. $U(s)=\sum_{j=1}^{d}\alpha_{u}^{(j)}h_{j}(s)$ which gives us $avar\left(\hat{\beta}_{-j,n}\right)=\sigma^{2}_{\epsilon}/{{\alpha_{x}^{(j)}}^{2}}$. Note that, excluding pathological cases, $\sigma^{2}_{U,j}=\int_{\mathcal{D}}(h_{j}(s)U(s))^{2}f_{S}(s)ds=\text{Var}(h_{j}(S)U(S))>0$ since $\alpha_{u}^{(j)}=0$. So the asymptotic variance of $\hat{\beta}_{j,n}^{PROJ}$ is strictly greater than $\sigma^{2}_{\epsilon}/{\alpha_{x}^{(j)}}^{2}$. This completes the proof of this case.
Case 2: $(\lambda_{d}=0)$: For this case we will use Lemma S3.6. Since the basis functions have been assumed to continuous and the domain $\mathcal{D}$ is compact, Assumption (i) of the lemma is immediately satisfied. Assumption (iii) is also satisfied for this basis function class for a choice of very small $\epsilon$ which corresponds a large enough $d$ since the unmeasured confounder is assumed to be a continuous function in the compact domain $\mathcal{D}$. Thus for $j\in O_{X,d}$, invoking the lemma we have $\|h_{j}U\|^{2}_{f_{S}}>\|h_{j}U_{>d}\|^{2}_{f_{S}}$. Few things to observe here is that $\|h_{j}U\|^{2}_{f_{S}}=\int_{\mathcal{D}}(h_{j}U)^{2}f_{S}ds=\sigma^{2}_{U,j}$. Next we see that $\sigma^{2}_{U,j,>d}=\int_{\mathcal{D}}(h_{j}U_{>d})^{2}f_{S}ds=\|h_{j}U_{>d}\|^{2}_{f_{S}}$. Thus, using this lemma, it can be directly inferred from Theorems 2 and 3 that $\sigma^{2}_{U,j}>\sigma^{2}_{U,j,>d}$ for $j\in O_{X,d}$. Thus even for $U$ having possibly infinite expansion i.e. for $\lambda_{d}=0$, if $\mathcal{H}$ satisfies the assumptions of Theorem 4 then:
This completes the proof of the theorem. ∎